OpenStudy (anonymous):

In a list of 100 numbers , the fraction of numbers greater than 15 is (19/25) and the fraction of numbers less than 15 is (7/50). How many numbers in the list are equal to 15

OpenStudy (texaschic101):

19/25(100) = 1900/25 = 76 (greater then 15) 7/50(100) = 700/50 = 14 (less then 15) 100 - (76 + 14) = 100 - 90 = 10 There are 10 numbers equal to 15.